# for Scalable MoE Pruning

### v1

@@ Abstract
Mixture-of-experts (MoEs) have been adopted for reducing inference costs by sparsely activating experts in Large language models (LLMs). Despite this reduction, the massive number of experts in MoEs still makes them expensive to serve. In this paper, we study how to address this, by pruning MoEs. Among pruning methodologies, unstructured pruning has been known to achieve the highest performance for a given pruning ratio, compared to structured pruning, since the latter imposes constraints on the sparsification structure. This is intuitive, as the solution space of unstructured pruning subsumes that of structured pruning. However, our counterintuitive finding reveals that expert pruning, a form of structured pruning, can actually precede unstructured pruning to outperform unstructured-only pruning. As existing expert pruning, requiring $O(\frac{k^{n}}{\sqrt{n}})$ forward passes for $n$ experts, cannot scale for recent MoEs, we propose a scalable alternative with $O(1)$ complexity, yet outperforming the more expensive methods. The key idea is leveraging a latent structure between experts, based on behavior similarity, such that the greedy decision of whether to prune closely captures the joint pruning effect. Ours is highly effective– For Snowflake Arctic, a 480B-sized MoE with 128 experts, our method needs only one H100 and two hours to achieve nearly no loss in performance with 40% sparsity, even in generative tasks such as GSM8K, where state-of-the-art unstructured pruning fails to. The code will be made publicly available.
@@ 1 Introduction
Large language models (LLMs) have become the state-of-the-art for various tasks OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Team et al. (2023); Lieber et al. (2024). However, their prohibitive inference cost is becoming a bottleneck to deployment Kaddour et al. (2023), and detrimental to the environment Strubell et al. (2019); Zeng et al. (2023).
@@
Despite the reduction in computational costs, the massive number of parameters remain unchanged, requiring significantly more GPU memory which makes serving large MoE models out of reach for many. To make matters worse, recent MoEs tend to increase the number of experts $n$, resulting in even larger MoEs. For instance, accommodating 56B parameters of Mixtral Jiang et al. (2024) with 8 experts or 132B of DBRX Databricks (2024) with 16 experts, or 480B of Snowflake Arctic Snowflake (2024) with 128 experts, requires an ever-increasing amount of memory and thus more GPUs to host and serve.
Figure 1: GSM8K 5-shot accuracy by pruning Snowflake Arctic, a 480B-sized MoE.
@@
To reduce the number of parameters, unstructured Frantar and Alistarh (2023); Sun et al. (2024), or structured pruning Ma et al. (2023) can be considered. Unstructured pruning allows weight tensors to be sparse anywhere. Structured pruning imposes a restricted pattern on sparsification, such as removing rows, entire weight tensors, or pruning experts in MoE Lu et al. (2024).
Unstructured pruning has been believed to achieve a higher pruning ratio, compared to structured pruning Chen et al. (2022), due to structural restriction. [^1] When comparing Table 1 in Ma et al. (2023) with Table 22 in Sun et al. (2024), the BoolQ performance drops from 71.22 to 61.44, by changing to structured pruning, even allowing less pruning ratio. This seems intuitive, as the solution space of unstructured pruning is a superset of the structured counterpart with structural restrictions. Consequently, an optimal solution that violates the restriction cannot be found from structured pruning, which may result in suboptimal outcomes.
In this paper, we show that Structured-Then-UNstructured pruning (STUN) outperforms unstructured pruning in MoEs. Though it may seem counter-intuitive, we demonstrate that a well-designed structured pruning, which is our expert pruning, would retain the performance after the pruning, and also ensure the pruned network remains robust to unstructured pruning that follows.
First, we argue that MoEs, by design, are ‘robust’ to expert pruning, as MoE training is intended to maintain performance after a number of experts have been pruned. This allows us to design expert pruning to maintain performance similar before and after pruning. To support this argument, we later draw a connection between standard MoE training and targeted dropout Gomez et al. (2019), a training method designed to enhance a network’s robustness to pruning.
Second, we argue that our proposed expert-pruned network remains robust to unstructured pruning. According to Mason-Williams and Dahlqvist (2024), higher kurtosis of weights, or having many outliers in the weight distribution deviating from the normal distribution, suggests more weights can be pruned while maintaining performance, or high robustness of unstructured pruning. We argue that, unlike unstructured pruning, expert (structured) pruning does not decrease kurtosis values, thereby maintaining the network’s robustness to unstructured pruning.
@@
However, existing expert-level pruning for MoE does not scale well over the solution space Lu et al. (2024), requiring an exhaustive combination of experts, leading to $O(\frac{k^{n}}{\sqrt{n}})$ GPU calls, with $k=\frac{1}{\phi^{\phi}(1-\phi)^{1-\phi}}$, and $\phi<1$ is pruning ratio. While this was acceptable in an early MoE work with few experts, it does not scale to recent trends in MoEs with large $n$ Bai et al. (2023); Dai et al. (2024); Snowflake (2024), or even infinity He (2024).
Our distinction is drastically reducing the complexity of expert pruning to $O(1)$, without compromising the performance. The main intuition is leveraging a latent structure between experts, based on behavior similarity, such that the greedy decision of whether to prune closely captures the joint pruning effect.
@@
  • We propose STUN, the first counterintuitive solution combining structured and unstructured pruning to break the pruning ratio limits of unstructured pruning, to the best of our knowledge.
  • We show that a well-designed expert pruning would retain the performance after the pruning, and the pruned network remains robust to unstructured pruning.
  • To materialize, we design $O(1)$ expert-level pruning method, which outperforms the previously proposed $O(\frac{k^{n}}{\sqrt{n}})$ solution Lu et al. (2024).
  • Our structured pruning, when followed by unstructured pruning, jointly optimizes pruning to achieve a state-of-the-art compression ratio. For Snowflake Arctic, a 480B-sized MoE with 128 experts, STUN needs only 1 H100 and two hours, achieving no loss up to 40% of sparsity, even for generative tasks such as GSM8K, where state-of-the-art unstructured pruning methods fail.
  • Code will be publicly available.
@@ 2.1 LLM Pruning
LLM pruning can be classified into unstructured and structured pruning Behnke and Heafield (2021). Unstructured pruning involves finding mask tensors to sparsify weight tensors. SparseGPT Frantar and Alistarh (2023) uses the Hessian matrix for second-order Taylor approximation, while GBLM-Pruner Das et al. (2024) and Pruner-Zero Dong et al. (2024) leverage gradients to identify mask tensors. However, as these methods demand substantial GPU memory for LLMs, we consider two recent works with higher memory efficiency and strong performance as our baselines: Wanda Sun et al. (2024) evaluates the importance of neurons in each layer by its weight multiplied by the activation value, removing those with low scores. While Wanda assumes a uniform pruning ratio across layers, OWL Yin et al. (2024) probes the optimal pruning ratio per layer, given the pruning budget.
Structured pruning, on the other hand, imposes constraints on the sparsification pattern, such as removing rows, columns, or even entire weight tensors. Early methods that involve pruning attention heads Voita et al. (2019); Shim et al. (2021); Zhang et al. (2021), rows Gong et al. (2022), entire dense layers Liang et al. (2021), or whole transformer blocks Fan et al. (2019); Li et al. (2020) fall under this category. Recent works have applied structured pruning for LLMs Ma et al. (2023); Cheng et al. (2024); Gao et al. (2024); Zhang et al. (2024); Dery et al. (2024), but without fine-tuning, these methods generally underperform compared to unstructured pruning.
Our distinction is to introduce a new class of pruning—structured-then-unstructured pruning—and demonstrate its significant advantages for MoEs, surpassing the performance of unstructured-only pruning. This approach differs from previous methods that combine structured and unstructured pruning Kurtic et al. (2022), which have not succeeded in outperforming unstructured pruning.
@@ 2.2 Expert Pruning
Early work on expert pruning was domain-specific, such as in translation MoEs, by keeping most activated experts Kim et al. (2021), or pruning based on gate statistics Koishekenov et al. (2023).
Lu et al. (2024) introduced a domain-agnostic expert pruning, searching for the best combination of experts to reduce the reconstruction loss, and quantify their criticality in output prediction.
Our distinction is eliminating the need for expensive combination enumeration, reducing the GPU calls from $O(\frac{k^{n}}{\sqrt{n}})$ to $O(1)$.
@@ 2.3 Pruning Robustness
Robustness in post-hoc pruning is quantified by whether performance is maintained after pruning. Training methods aimed at enhancing this robustness include applying group lasso Wen et al. (2016); Behnke and Heafield (2021), L1 Han et al. (2015); Liu et al. (2017), or L0 Louizos et al. (2018) regularization to the weights, as well as learning a sparse scaling factor Huang and Wang (2018); Li et al. (2020); Gong et al. (2022). Additionally, Targeted Dropout Gomez et al. (2019) demonstrates that applying stochastic dropout to less important parts of the network can increase robustness to pruning.
Meanwhile, kurtosis of weights Mason-Williams and Dahlqvist (2024) has been used as a proxy of robustness, stating networks with higher weight kurtosis can tolerate higher unstructured pruning ratios.
Our contribution lies in connecting these findings for designing expert pruning. We argue that MoE training inherently enhances robustness to expert-level pruning, and this coarse structured pruning subsequently makes the remaining network more resilient to unstructured pruning. In section 5, we argue the former, based on the resemblance of MoE training and targeted dropout, and the latter, inspired by Mason-Williams and Dahlqvist (2024).
@@ 3 Preliminaries: MoE
4 Structured-Then-UNstructured Pruning (STUN)
4.1 Overview
Essentially, STUN first performs structured (expert) pruning, until the loss is negligible, then followed by unstructured pruning, such as OWL Yin et al. (2024) or Wanda Sun et al. (2024).
Our key contribution is to replace combinatorial loss with $O(1)$ expert pruning method, by leveraging latent cluster structure among experts, based on behavioral similarity.
Specifically, we find clusters of similar experts layer by layer, yielding a total of $\phi nl$ clusters in the whole MoE, where $\phi$ is the pruning ratio, $n$ is the number of experts in each layer, and $l$ is the number of layers in MoE. Then we greedily prune every expert but one representative per each cluster.
Later sections show why our greedy pruning is as effective as its combinatorial counterpart, and how it effectively connects to MoE training beforehand, and unstructured pruning afterwards.
4.2 $O(\frac{k^{n}}{\sqrt{n}})$: Combinatorial Reconstruction Loss
Now we will derive our algorithm in section 4.1 formally, starting from the conventional goal of pruning– minimizing the reconstruction loss. Reconstruction loss has been employed to assess how closely the pruned model $\theta-\theta_{S}$ without expert set $S$ mirrors the behavior of the unpruned $\theta$  Lu et al. (2024). Formally, this loss is quantified by the Frobenius norm of the difference between the original output $M(x;\theta)$ and the output of pruned layer $M(x;\theta-\theta_{S})$, denoted as $\mathcal{E}_{S}$.
@@ 4.2 $O(\frac{k^{n}}{\sqrt{n}})$: Combinatorial Reconstruction Loss
To elaborate, deciding which experts to prune using Eq. 4 for $|S|=\phi n$ requires (${n\choose{|S|}}\approx O(\frac{k^{n}}{\sqrt{n}})$ forward passes according to Stirling’s approximation, where $k=\frac{1}{\phi^{\phi}(1-\phi)^{1-\phi}}$, and $\phi<1$ represents the pruning ratio. [^2] In detail, 23951146041928082866135587776380551750 forward passes per layer at minimum, for massive MoEs with $n=128$ experts such as Snowflake Arctic Snowflake (2024). Our distinction is to lower the computation to $O(1)$, without compromising the performance– In fact, as we will elaborate later, we outperform the combinatorial objective.
4.3 Towards $O(n)$: Probabilistic Interpretation
As a stepping stone towards $O(1)$, we propose to rephrase the goal of finding $\theta_{S}$ to minimize $\mathcal{E}_{S}$ as:
$\textrm{argmax}_{S}P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})$ (5)
where $s_{i}$s are the experts included in the expert set $S$, and $P(X_{1}=s_{1},\cdots,X_{|S|}=s_{|S|})$ is the joint probability of pruning $S$.
Section 4.2 corresponds to enumerating joint probability from all combinations, requiring ($n\choose{|S|}$ different values, which is compute-intensive. When chain rule is applied, Eq. 5 can be reformulated as follows:
@@ 4.3 Towards $O(n)$: Probabilistic Interpretation
We propose a greedy optimization for Eq. 6– We decompose the multiplication of Eq. 6 and at each step $k$, and obtain the distribution $P(X_{k}|s_{1},\cdots,s_{k-1})$, to select $X_{k}$ that maximizes the probability. For simplicity, we will omit $X_{k}$s from this point on.

### v2

@@ Abstract
Mixture-of-experts (MoEs) have been adopted to reduce inference costs by sparsely activating experts in large language models (LLMs). Despite these reductions, the massive number of parameters in MoEs still makes them expensive to serve. Conventionally, unstructured or structured pruning has been considered to reduce the number of parameters. Our key contribution is exploring the interpolation between structured and unstructured pruning, to propose a novel structured-then-unstructured (STUN) approach outperforming both structured and unstructured pruning, especially for MoEs. In the first stage, we show a scalable expert pruning with O(1) forward pass, unlike existing work requiring $O(\frac{k^{n}}{\sqrt{n}})$ forward passes for $n$ experts that cannot scale for recent MoEs with hundreds of experts. We then show our expert-pruned MoEs are robust to unstructured pruning to follow. Experiments on Snowflake Arctic and Mixtral show that our proposal is highly effective– For Snowflake Arctic, a 480B-sized MoE with 128 experts, our method needs only one H100 and two hours to achieve nearly no loss in performance with 40% sparsity, even in generative tasks such as GSM8K, where state-of-the-art structured or unstructured pruning methods fail. The code is publicly available. [^1] https://github.com/thnkinbtfly/STUN
@@ 1 Introduction
Large language models (LLMs) have become state-of-the-art for various tasks OpenAI (2023); Touvron et al. (2023); Jiang et al. (2023); Lieber et al. (2024). However, their prohibitive inference cost is becoming a bottleneck to deployment Kaddour et al. (2023), and detrimental to the environment Strubell et al. (2019); Zeng et al. (2023).
@@
Despite these reductions, the massive number of parameters remains unchanged, requiring significantly more GPU memory, which makes serving large MoE models challenging for many. Additionally, recent MoEs tend to increase the number of experts $n$, resulting in even larger MoEs. For instance, accommodating 56B parameters of Mixtral Jiang et al. (2024) with 8 experts or 132B of DBRX Databricks (2024) with 16 experts, or 480B of Snowflake Arctic Snowflake (2024) with 128 experts, requires an ever-growing amount of memory and more GPUs to serve.
Figure 1: GSM8K 5-shot accuracy by pruning Mixtral-8x7B-Instruct by 50% of sparsity. We probe interpolation of structured and unstructured pruning, by varying the ratio of structured pruning.
@@
To reduce the number of parameters, unstructured Frantar and Alistarh (2023); Sun et al. (2024), or structured pruning Ma et al. (2023) can be considered. Unstructured pruning allows weight tensors to be sparse anywhere, while structured pruning imposes patterns on sparsification, such as removing rows, entire weight tensors Ma et al. (2023), or pruning experts in MoE Lu et al. (2024a).
In this paper, we propose the interpolation of two, Structured-Then-UNstructured pruning (STUN). Figure 1 motivates our interpolated method, where unstructured- or structured-only, $x=0$ and $x=1$, respectively, is outperformed by the peak, combining both.
@@
Our first phase, expert pruning, by leveraging model-inherent expert for pruning, significantly outperforms row/column-level structured pruning (Figure 1 blue; Ma et al., 2023) However, existing expert-level pruning for MoE (Figure 1 grey; Lu et al., 2024a) often does not scale well over the solution space, requiring an exhaustive combination of experts, leading to $O(\frac{k^{n}}{\sqrt{n}})$ GPU calls, with $k=\frac{1}{\phi^{\phi}(1-\phi)^{1-\phi}}$, and $\phi<1$ is sparsity Lu et al. (2024a). While this was acceptable in an early MoE work with few experts, it does not scale to recent trends in MoEs with large $n$ Bai et al. (2023); Dai et al. (2024); Snowflake (2024), or even infinity He (2024). Our distinction is drastically reducing the number of GPU calls to $O(1)$, without compromising the performance. The main intuition is leveraging a latent structure between experts, based on behavior similarity, such that the greedy decision of whether to prune closely captures the joint pruning effect.
The second contribution is allowing unstructured phase to follow, to consider both inter- and intra-expert sparsity (Table 1). STUN removes redundant experts by expert-level structured pruning first, then desires fine-grained sparsity within individual experts. [^2] From now on, we will define sparsity as the number of pruned parameters divided by the total number of parameters in the original model.
We support STUN with the findings of Mason-Williams and Dahlqvist (2024), which show that higher kurtosis in the weight distribution (indicating many outliers) suggests more weights can be pruned while maintaining performance, highlighting the robustness of unstructured pruning. We argue that expert-level pruning does not reduce kurtosis, thereby preserving the network’s resilience to unstructured pruning.
Inter-Expert Sparsity Intra-Expert Sparsity Figure 1
Lu et al., 2024a ($\bullet$)
LLM-Pruner ($\bullet$), Wanda, OWL ($\bullet$)
Ours ($\bullet$)
Table 1: Comparison of existing pruning methods for MoEs.
@@
  • We propose STUN, the first method to combine structured and unstructured pruning, outperforming both approaches.
  • Scalable first phase: We design an expert-level pruning method with $O(1)$ GPU calls, outperforming the $O(\frac{k^{n}}{\sqrt{n}})$ solution Lu et al. (2024a).
  • Justifying the second phase: We show the expert-pruned network remains robust to unstructured pruning to follow.
  • State-of-the-art efficiency and compression: For Snowflake Arctic (480B, 128 experts), it requires just 1 H100 and two hours, with no backpropagation or fine-tuning needed. Compression reaches up to 40% sparsity without compromising performance, even in generative tasks like GSM8K, where unstructured pruning fails. We report consistent results for Mixtral models.
@@ 2.1 LLM Pruning
LLM pruning can be classified into unstructured and structured pruning Behnke and Heafield (2021). Unstructured pruning involves finding mask tensors to sparsify weight tensors. Such masking leads to practical speedups in hardware such as CPU NeuralMagic (2021), and ongoing research is actively developing methods to achieve similar speedups on GPUs Mishra et al. (2021); Zhao et al. (2024). SparseGPT Frantar and Alistarh (2023) uses the Hessian matrix for second-order Taylor approximation, while GBLM-Pruner Das et al. (2024) and Pruner-Zero Dong et al. (2024) leverage gradients to identify mask tensors. However, as these methods demand substantial GPU memory for LLMs, we focus on more memory-efficient approaches, using two recent baselines: Wanda Sun et al. (2024) evaluates the importance of neurons in each layer by its weight multiplied by the activation value, removing those with low scores. While Wanda assumes a uniform sparsity across layers, OWL Yin et al. (2024) probes the optimal sparsity per layer, given the pruning budget.
Structured pruning, on the other hand, imposes constraints on the sparsification pattern, such as removing rows, columns, or even entire weight tensors. Early methods that involve pruning attention heads Voita et al. (2019); Shim et al. (2021); Zhang et al. (2021), rows Gong et al. (2022), entire dense layers Liang et al. (2021), or whole transformer blocks Fan et al. (2019); Li et al. (2020) fall under this category. Recent works have applied structured pruning for LLMs Ma et al. (2023); Cheng et al. (2024); Gao et al. (2024); Zhang et al. (2024a); Dery et al. (2024), but without fine-tuning, these methods generally underperform when compared to unstructured pruning.
Our distinction is to introduce a new class of pruning– structured-then-unstructured pruning– and demonstrate its significant advantages for MoEs, surpassing the performance of either method alone. This approach differs from previous methods that combine structured and unstructured pruning Kurtic et al. (2022), which failed to outperform unstructured pruning.
@@ 2.2 Expert Pruning
Early work on expert pruning was domain-specific Kim et al. (2021); Koishekenov et al. (2023); Liu et al. (2024), such as in translation MoEs, by keeping most activated experts Kim et al. (2021), or pruning based on gate statistics Koishekenov et al. (2023). Lu et al. (2024a) introduced a domain-agnostic expert pruning, searching for the best combination of experts to reduce the reconstruction loss, and quantify their criticality in output prediction. Concurrently to our study, Zhang et al. (2024b) proposed an efficient expert pruning method.
Our distinction is two-fold. First, we interpolate expert pruning with unstructured pruning to outperform either method alone. Second, for scalable expert-level structured pruning, we derive a scalable expert-level structured pruning method with $O(1)$ GPU calls, improving on the $O(\frac{k^{n}}{\sqrt{n}})$ solution enumerating combinatorial pruning.
@@ 2.3 Pruning Robustness
Robustness in post-hoc pruning is quantified by whether performance is maintained after pruning. Kurtosis of weights Mason-Williams and Dahlqvist (2024) has been used as a proxy for robustness, with networks showing higher weight kurtosis able to tolerate higher unstructured pruning ratios. Our contribution is demonstrating that an expert-pruned network remains robust to additional unstructured pruning, which naturally supports our design of unstructured pruning as the second phase.
@@ 3 Preliminaries: MoE
4 Proposed Method
Figure 2: Overview of our proposed STUN. ➀ We first remove redundant experts with expert-level structured pruning, then ➁ perform unstructured pruning inside individual experts. Black box represents a layer in MoE, and different colors represent different behavioral similarities.
4.1 Overview: STUN
Figure 2 overviews our two-phase approach, interpolating structured and unstructured pruning as motivated in Table 1. Section 4.2 describes ➀ how we remove redundant experts with expert-level structured pruning with high scalability, then Section 4.3 describes ➁ how we perform unstructured pruning inside individual experts.
4.2 Expert-level Structured Pruning with $O(1)$ GPU calls
Now, we describe our expert-level structured pruning with $O(1)$ GPU calls. [^3] We focus on GPU cost, since it dominates the CPU cost– For example, the accumulated CPU cost (including the hyperparameter search to meet desired sparsity) required by our algorithm is less than 1 minute, even on 480B Snowflake Arctic. Previous solution Lu et al. (2024a) minimizes reconstruction loss (Section 4.2.1), requiring GPU call per combination of experts, that is $O(\frac{k^{n}}{\sqrt{n}})$. Our key contribution is to approximate this combinatorial reconstruction loss to reduce the number of GPU calls to $O(1)$, by leveraging latent cluster structure among experts, based on behavioral similarity.
Specifically, we find clusters of similar experts layer by layer, yielding a total of $\phi nl$ clusters in the whole MoE, where $\phi$ is the sparsity, $n$ is the number of experts in each layer, and $l$ is the number of layers in MoE. Then we greedily prune every expert but one representative per each cluster.
Later sections show why our greedy pruning is as effective as its combinatorial counterpart.
4.2.1 $O(\frac{k^{n}}{\sqrt{n}})$: Combinatorial Reconstruction Loss
We start from the conventional goal of pruning– minimizing the reconstruction loss. Reconstruction loss has been employed to assess how closely the pruned model $\theta-\theta_{S}$ without expert set $S$ mirrors the behavior of the unpruned $\theta$  Lu et al. (2024a). Formally, this loss is quantified by the Frobenius norm of the difference between the original output $M(x;\theta)$ and the output of pruned layer $M(x;\theta-\theta_{S})$, denoted as $\mathcal{E}_{S}$.
@@ 4.2.1 $O(\frac{k^{n}}{\sqrt{n}})$: Combinatorial Reconstruction Loss
To elaborate, deciding which experts to prune using Eq. 4 for $|S|=\phi n$ requires (${n\choose{|S|}}\approx O(\frac{k^{n}}{\sqrt{n}})$ forward passes according to Stirling’s approximation, where $k=\frac{1}{\phi^{\phi}(1-\phi)^{1-\phi}}$, and $\phi<1$ represents the sparsity. Our distinction is to lower the computation to $O(1)$, without compromising the performance– In fact, as we will elaborate later, we outperform the combinatorial objective.
4.2.2 Towards $O(n)$: Probabilistic Interpretation
As a stepping stone towards $O(1)$, we propose to rephrase the goal of finding $\theta_{S}$ to minimize $\mathcal{E}_{S}$ (Eq. 4) as:
@@ 4.2.2 Towards $O(n)$: Probabilistic Interpretation
Our contribution is greedy optimization without compromise for Eq. 5– We decompose the multiplication of Eq. 5 at each step $k$, and obtain the distribution $P(X_{k}|s_{1},\cdots,s_{k-1})$, to select $X_{k}$ that maximizes the probability. To achieve it, we estimate the rank between the probabilities. Such rank estimation can benefit from the latent structure among experts, specifically, a cluster of similar experts in MoE. Given cluster mapping $c$ which maps an expert to a set of similarly behaving experts, we assign the value $P(E_{i}|S_{k-1})$, as follows:
@@
This enables the calculation of all $P(E_{i}|S_{k-1})$ in Eq. 5 from $P(E_{i})$s, which needs only $n$ forwards in total.
@@
Our remaining task is to obtain cluster information $c$: One signal is pairwise behavioral similarity $b_{i,j}$, from the pretrained router weights $W$ at a minimal cost. Suppose two rows $W_{i}\approx W_{j}$ are similar; then $r_{i}(x)\approx r_{j}(x)$, meaning $E_{i},E_{j}$ tend to be selected by similar inputs, implying similar expertise. Thus, the behavioral similarity $b_{i,j}$ between two experts $E_{i},E_{j}$ can be obtained as follows: